# Future-Oriented Navigation: Dynamic Obstacle Avoidance with One-Shot Energy-Based Multimodal Motion Prediction

### v1

@@ I INTRODUCTION
Factory, warehouse, and laboratory layouts are subject to frequent changes, and the presence of humans and manually driven vehicles further complicates motion planning for Autonomous Mobile Robots (AMRs). To ensure collision avoidance, AMRs must anticipate interactions with static and dynamic objects, though the unpredictability of humans poses significant challenges. A prevalent industrial strategy is to halt the robot when obstacles are detected [1], risking downtime if obstructions persist or deadlocks occur. This study addresses collision-free navigation for AMRs in industrial settings but can be generalized to other settings with similar traffic dynamics.
Dynamic Obstacle Avoidance (DOA) has been a key research area since the advent of AMRs, with an initial emphasis on safety. This often leads to overly cautious behavior, exemplified by the Freezing Robot Problem (FRP) [2], where the planner deems all forward paths and stops the robot. Advancements in perception and computation technologies enable robots to rapidly and accurately gather extensive environmental data and coordinate their actions, improving their ability to implement more effective decision-making strategies. Recent research has improved motion planning efficiency and addressed FRP by incorporating human movements and intentions [3, 4], especially in social robot navigation [5]. In traditional model-based approaches, the Constant Velocity Model (CVM) and velocity-based strategies are widely used for predicting human motion in navigation frameworks [3, 6, 7]. Additionally, predefined kinematics-based [8] and interaction-based [9] models offer more detailed insights for motion planners. Recently, learning-based models have become the primary motion predictor in most prescient navigation frameworks due to their high flexibility and ability to interpret complex and uncertain interactions and environments [9, 4, 10, 11].
@@
This work presents a navigation framework for multiple AMRs in dynamic environments, tailored for industrial transport tasks, incorporating multimodal motion prediction, prediction grouping, static obstacle avoidance, proactive DOA, and fleet collision avoidance. The approach integrates a data-driven motion predictor with a predictive trajectory planner. Specifically, an energy-based model predicts possible future motions of dynamic obstacles as probability maps, which are then used by distributed Model Predictive Control (MPC) to ensure safe navigation. Multiple tests are designed to evaluate the method’s performance in a comparative study with state-of-the-art approaches. The main contributions of our work are threefold:
  • Introducing a stable energy-based learning approach for multimodal motion prediction enabling improved integration with downstream planning and control.
@@
  • Evaluating the framework under various conditions and scenarios, with and without cooperative pedestrians, and validating multiple design choices.
@@ VII-A Motion Prediction Evaluation
In pure motion prediction tasks, the capture of diversity is emphasized to increase the likelihood of matching the ground-truth trajectory. In the context of downstream planning, the focus is on ensuring the predictions are closely aligned with the ground-truth probability distribution. As implied in Fig. 2, training via NLL produces concentrated predictions. Since real-world datasets don’t come with ground-truth distributions and it is difficult to estimate a distribution due to the lack of data, for quantitative analysis, a simulated crossroad scene is used as shown in Fig. 3, where a pedestrian may go straight, turn left or right at the intersection. To increase local uncertainty, the right side is an open area, and pedestrians randomly pick endpoints along the right edge of the map. The ground-truth distribution is estimated by fitting samples into a GMM (bottom left corner in Fig. 3) and compared to output probability maps from different models in terms of Jensen-Shannon Divergence (JSD) and Earth Mover’s Distance (EMD, or the Wasserstein distance [17]). JSD measures overall similarity rather than the actual distance between distribution peaks. EMD calculates the minimum “effort” to transform one distribution into another, which is the product of the amount of moved probability mass and the corresponding distance. EMD depends on both the distance between distributions and their overall shapes. Given EMD’s computational complexity in high dimensions, Sliced Wasserstein Distance (SWD) [31] is chosen for Monte Carlo approximation. As shown in TABLE I, KLD and NLL perform similarly on JSD, while NLL outperforms others on SWD. In our experiments, the original NLL loss performs adequately in the simple crossroad dataset. However, without the PELU and soft loss modifications, the training becomes unstable and prone to failure for real-world and complex datasets. The model also generates severe noise without the PELU layer, making it infeasible for downstream tasks.
@@ VII-C Long-term Robustness
Apart from Scenario 4, to further examine the robustness of the proposed pipeline, more long-term running experiments in the warehouse (with cooperative SF pedestrians) are implemented. To increase the uncertainty and interaction frequency, three AMRs and three pedestrians operating in close proximity are included, set to follow predefined routes. The location and timing of interactions between AMRs and pedestrians varied randomly due to the human motion noise. In three trials, each AMR encountered over 20 pedestrian interactions (defined as instances where a pedestrian influenced the robot’s actions) and multiple fleet collision avoidance cases. The proposed approach demonstrates robustness and adaptability throughout the tests. Robots can autonomously select different strategies based on the situation, such as taking detours when necessary, cutting shorter paths if safe, reversing and restarting in response to dynamic obstacles, etc. By leveraging distributed MPC, collision avoidance between multiple AMRs is also handled smoothly.
@@ VII-D Simulation in Gazebo
The proposed pipeline is deployed in Gazebo to achieve a realistic simulation. Fig. 7 demonstrates the warehouse and hospital environments. The corresponding code and environments are available in the provided repository.
@@
VII-E Ablation Study of the Normalized Mask
In different loss functions, different forms of ground truth are employed, typically either the estimated ground-truth probability map $\bm{A}$ or its normalized version $\bar{\bm{A}}$. To understand the impact of ground-truth formulation on training outcomes, we further investigated its role across various losses. For the BCE loss, since it requires a Sigmoid output layer, the target label must be between 0 and 1 for mathematical soundness. If the mask is not normalized, the BCE loss becomes inapplicable. For the KLD loss, as defined in (15), normalizing the ground-truth mask scales the loss by the reciprocal of the maximum value in $\bm{A}$. When evaluating the effect of normalization using the same setup as in TABLE I, the JSD of KLD with normalized mask is $0.3267$ and the SWD is $9.6767\times 10^{-5}$, which are slightly better than KLD without normalized mask (JSD: $0.3407$, SWD: $9.958^{-5}$) but not significantly. Moreover, normalization had no essential impact on obstacle avoidance performance as it cannot solve the problem of over-conservative predictions. For the NLL and ENLL loss, using the normalized mask instead of the raw probability map scales up the negative component of the loss, which may bias training toward stronger alignment with the mask. In practical tests across the warehouse and hospital scenarios, ENLL trained with unnormalized masks produced slightly more concentrated predictions, but this did not lead to significant differences in obstacle avoidance performance.
@@ VIII Conclusion and Future Work
One limitation of the current work is that tests are conducted in simulated environments and the complex combinations of static and dynamic obstacles haven’t been further studied, which is the primary focus of our future research. Other future work could include three aspects: extending the pipeline to include the vision and tracking system; implementing a local reference generator for MPC to achieve fast convergence; and considering more realistic transport tasks involving more mobile robots and a task scheduler.

### v2

@@ Abstract
Index Terms:
Human-aware motion planning, collision avoidance, deep learning methods.
@@ I Introduction
Factory, warehouse, and laboratory layouts are subject to changes, and the presence of humans and manually driven vehicles further complicates motion planning for Autonomous Mobile Robots (AMRs). To ensure collision avoidance, AMRs must anticipate interactions with static and dynamic objects, though the unpredictability of humans poses significant challenges. A prevalent strategy is to halt the robot when obstacles are detected [1], risking downtime if obstructions persist or deadlocks occur. This study addresses collision-free navigation for AMRs in industrial settings but can be generalized to other settings with similar traffic dynamics.
Dynamic Obstacle Avoidance (DOA) has been a key research area since the advent of AMRs, with an initial emphasis on safety. This often leads to overly cautious behavior, exemplified by the Freezing Robot Problem (FRP) [2], where the planner deems all paths and stops the robot. Advancements in perception and computation technologies enable robots to rapidly and accurately gather extensive environmental data and coordinate actions, improving the ability to implement effective decision-making strategies. Recent research has improved motion planning efficiency and addressed FRP by incorporating human movements and intentions [3, 4], especially in social robot navigation [5]. In traditional model-based approaches, the Constant Velocity Model (CVM) and velocity-based strategies are widely used for predicting human motion in navigation frameworks [3, 6, 7]. Additionally, predefined kinematics-based [8] and interaction-based [9] models offer more detailed insights for motion planners. Recently, learning-based models have become the primary motion predictor in most prescient navigation frameworks due to their high flexibility and ability to interpret complex and uncertain interactions and environments [9, 4, 10, 11].
@@
This work presents a navigation framework for multiple AMRs in dynamic environments, tailored for industrial transport tasks, incorporating multimodal motion prediction, prediction grouping, static obstacle avoidance, proactive DOA, and fleet collision avoidance. The approach integrates a data-driven motion predictor with a predictive trajectory planner. Specifically, an energy-based model predicts possible future motions of dynamic obstacles as probability maps, which are used by distributed Model Predictive Control (MPC) for safe navigation. Multiple tests are designed to evaluate the method’s performance in a comparative study with state-of-the-art approaches. The main contributions are threefold:
  • Introducing a fast and stable energy-based learning approach for multimodal motion prediction enabling improved integration with downstream planning and control.
@@
  • Evaluating the framework under various conditions and scenarios, with and without cooperative pedestrians.
@@ VII-A Motion Prediction Evaluation
In motion prediction tasks, the capture of diversity is emphasized to increase the likelihood of matching the ground-truth trajectory. For downstream planning, the focus is on ensuring the predictions are closely aligned with the ground-truth probability distribution. As implied in Fig. 2, training via NLL produces concentrated predictions. Since real-world datasets don’t come with ground-truth distributions and it is difficult to estimate a distribution due to the lack of data, for quantitative analysis, a simulated crossroad scene is used as shown in Fig. 3, where a pedestrian may go straight, turn left or right at the intersection. To increase local uncertainty, the right side is an open area, and pedestrians randomly pick endpoints along the right edge of the map. The ground-truth distribution is estimated by fitting samples into a GMM (bottom left corner in Fig. 3) and compared to output probability maps from different models in terms of Jensen-Shannon Divergence (JSD) and Earth Mover’s Distance (EMD, or the Wasserstein distance [17]). JSD measures overall similarity rather than the actual distance between distribution peaks. EMD calculates the minimum “effort” to transform one distribution into another, which is the product of the amount of moved probability mass and the corresponding distance. EMD depends on both the distance between distributions and their overall shapes. Given EMD’s computational complexity in high dimensions, Sliced Wasserstein Distance (SWD) [31] is chosen for Monte Carlo approximation. As shown in TABLE I, KLD and NLL perform similarly on JSD, while NLL outperforms others on SWD. In our experiments, the original NLL loss performs adequately in the simple crossroad dataset. However, without the PELU and soft loss modifications, the training is unstable and prone to failure for real-world and complex datasets. The model also generates severe noise without the PELU layer, making it infeasible for downstream tasks.
@@ VII-C Long-term Robustness
Apart from Scenario 4, to further examine the robustness of the proposed pipeline, long-term running experiments in the warehouse (with cooperative SF pedestrians) are implemented. To increase the uncertainty and interaction frequency, three AMRs and three pedestrians operating in close proximity are included, set to follow predefined routes. The location and timing of interactions between AMRs and pedestrians varied randomly due to the human motion noise. In three trials, each AMR encountered over 20 pedestrian interactions (defined as instances where a pedestrian influenced the robot’s actions) and multiple fleet collision avoidance cases. The proposed approach demonstrates robustness and adaptability throughout the tests. Robots autonomously select different strategies based on the situation, such as taking detours when necessary, cutting shorter paths if safe, reversing in response to dynamic obstacles, etc. With distributed MPC, collision avoidance between multiple AMRs is also handled smoothly.
@@ VII-D Simulation in Gazebo
The proposed pipeline is deployed in Gazebo for realistic simulation. Fig. 7 demonstrates the warehouse and hospital environments. The corresponding code and environments are available in the provided repository.
@@
VII-E Relevant Hyperparameters and Normalized Masks
In this session, we briefly discuss some relevant hyperparameters and analyze the use of normalized masks. In the original Gibbs-Boltzmann distribution, there is an inverse temperature parameter scaling energy values in the numerator of (12), which is normally 1 as in this work. In practice, this can be used to tune the conservativeness of the prediction. Another parameter that affects the conservatives of outputs is the variance of the ground-truth Gaussian mask $\bm{A}$. BCE and KLD are more sensitive to the variance of the mask, as they attempt to match it exactly. The proposed ENLL loss is less affected by the mask width, offering greater robustness.
In different loss functions, different forms of masks are employed, either the estimated ground-truth probability map $\bm{A}$ or its normalization $\bar{\bm{A}}$. For the BCE loss, normalization is necessary for mathematical soundness since it requires a Sigmoid output layer. Otherwise, it becomes inapplicable. For the KLD loss, as defined in (15), normalizing the mask scales it by the reciprocal of the maximum value in $\bm{A}$. When evaluating using the setup in TABLE I, with normalized masks, the JSD is $0.3267$ and the SWD is $9.6767\times 10^{-5}$, which are slightly better than KLD without normalized masks but not significantly. Moreover, normalization had no essential impact on obstacle avoidance performance as it cannot solve the problem of over-conservativeness. For the ENLL loss, using the normalized mask scales up its negative component, which may bias training toward stronger alignment with the mask. In practice, ENLL trained with unnormalized masks produced slightly more concentrated predictions but did not lead to significant differences in obstacle avoidance performance.
@@ VIII Conclusion and Future Work
One limitation of this work is the use of simulated environments, with complex static–dynamic obstacle interactions left for future investigation. Other future work could include three aspects: extending the pipeline to include the vision and tracking system; implementing a reference generator for MPC to achieve fast convergence; and considering industrial transport tasks involving more mobile robots and task schedulers.